darkknight:

Precalc: Its not necessarly a question but I kind of need review on this concept. So we learned about linear rational functions. Rational functions are in the form of f(x)= p(x)/q(x) where p(x) and q(x) are both polynomial. We also learned how to find horizontal and vertical asymptotes and I'm fine with that. So to write an equation in standard form I know how to do it systems of eqns (ax+b)/(x+c)=y then solve for each variable. But I need a little review. Can someone post a problem that has either 3 points, 2 points and 1 asymptote or 2 asymptotes and 1 point

darkknight:

k, ill just do a problem and someone check my work: points (0,3) (4,2) and (7,6) (a0+b)/(0+c)=3 >> 3c=b (a4+b)/(4+c) =2 >> 8+2c =4a +b >> 4a+b-2c=8 (a7+b)/(7+c)=6 >> 42+6c = 7a+b >> 7a+b-6c=42 7(4a+b-2c=8) >> 28a+7b-14c = 56 4(7a+b-6c = 42) >> 28a+4b-24c = 168 28a+4b-24c=168)-(28a+7b-14c=56) >> -3b-10c= 112 -3(3c)-10c=112 >> -19c = 112 >> c = -112/19 3(-112/19) = b = -336/19 4a+(-336/19) - (2(-112/19))=8 8(19/19)= 152/19 4a - (-224/19)= 488/19 4a +224/19 = 488/19 4a = 264/19 a = 66/19 ((66/19)x+(-336/19))/((x)+(-112/19))?

justjm:

I'll check your thing later when I get a chance..but basically are you trying to look for practice? Perhaps check KhanAcademy, or some other online practice carrier. Does your teacher give you homework? Maybe delve in more into your textbook? Do you want a few practice materials I had when I took precalc?

darkknight:

sure

justjm:

If that is for 3 points, then you are correct